OpenStudy (anonymous):

what is the solution of the proportion 8/x+2 = 3/x-1?

OpenStudy (anonymous):

\[\frac{ 8 }{ x+2 }=\frac{3 }{ x-1 }\] 8(x-1)=3(x+2) 8x-8=3x+6 8x-3x=6+8 5x=14, \[x=\frac{ 14 }{ 5 }=2.8\]